OpenStudy (anonymous):

Expand the Expression x>0 , y>0 , z>0 log (x^2/yz^4)

OpenStudy (amistre64):

log a/b = loga - logb log ab = loga + logb

OpenStudy (amistre64):

log n^r = r logn

OpenStudy (anonymous):

him can you elaborate more please

OpenStudy (amistre64):

not really, those are the basic rules for log operations. I can complicate them, but I doubt I can make them any easier ...

OpenStudy (anonymous):

can you giveme an example and work it out

OpenStudy (amistre64):

ok, but its just gonna be the same thing but with numbers in there

OpenStudy (amistre64):

log (1/2) = log(1)-log(2) log(4*3) = log(4)+log(3) log(6^3) = 3 log(6)

OpenStudy (amistre64):

youll need all these operations to expand the one youve got

OpenStudy (anonymous):

so the final answer would be 3log(6)?

OpenStudy (anonymous):

loga (x^2/yz^4)--what do i do with a

OpenStudy (amistre64):

"a" stays the same unless your asked to change of base it

OpenStudy (amistre64):

log (x^2/yz^4) do you see the division sign? the fraction bar? split this into its subtraction parts

OpenStudy (anonymous):

yesI understand it now! thanks

OpenStudy (amistre64):

log (x^2/yz^4) = log (x^2) - log (yz^4) where you see the multiplication, split it into its addition parts log (x^2/yz^4) = log (x^2) - log (y)+ log(z^4) and where you see exponents, turn it into its like parts

OpenStudy (amistre64):

log (x^2/yz^4) = 2log (x) - log (y)+ 4log(z)

OpenStudy (amistre64):

the base "a" doesnt change thruout it unless directed otherswise

OpenStudy (amistre64):

i might have a typo in there :)

OpenStudy (amistre64):

log (x^2/yz^4) = log (x^2) - (log (y)+ log(z^4)) is better since the hole of it is subtracted log (x^2/yz^4) = log (x^2) - log (y) - log(z^4) and then its the same from there